OpenStudy (anonymous):

Solve 3x^2 = -30x + 24. ( quadratic formula)

OpenStudy (loser66):

what is the "quadratic formula answer."?

OpenStudy (anonymous):

It says that you have to use the quadratic formula to solve it

OpenStudy (loser66):

do you know that formula?

OpenStudy (anonymous):

http://www.purplemath.com/modules/quadform.htmhttp://www.purplemath.com/modules/quadform.htm I just don't get it completely

OpenStudy (uri):

First arrange the equation.. 3x^2 = -30x + 24. 30x^2+30x-24=0 Now factorize with 3 3(x^2+10x-8=0)

OpenStudy (uri):